What amount would parents of a 5-year-old boy need to deposit into an account today, so that they will have $150,000 for their child’s college tuition 14 years from now? The money will receive 8.1%, compounded every-other-month? Round off to 2 decimal places.

Answer by nerdybill(7384)

What amount would parents of a 5-year-old boy need to deposit into an account today, so that they will have $150,000 for their child’s college tuition 14 years from now? The money will receive 8.1%, compounded every-other-month? Round off to 2 decimal places.
.
A = P(1+i/6)^(6*n)
150000 = P(1+.081/6)^(6*14)
150000 = P(1.0135)^(6*14)
150000 = P(1.0135)^(84)
150000 = P(3.085)
150000/3.085 = P
$48629.07 = P